1. Cost figures for a hypothetical firm are given in the following table. Use them for the exercises below. The firm is selling in a perfectly competitive market.
OUTPUT FIXED COST AFC VARIABLE COST AVC TOTAL COST ATC MC
1 $50 $30
2 $50 $50
3 $50 $80
4 $50 $120
5 $50 $170

a. Fill in the blank columns.
b. What is the minimum price needed by the firm to break even?
c. What is the shutdown price?
d. At a price of $40, what output level would the firm produce? What would its profits be?
1
Expert's answer
2020-07-02T18:06:15-0400

a)



b)


"Q=\frac {FC}{p-AVC}"


"3=\frac {50}{p-27}"


"p=43.7"

c)


"5=\frac {50}{p-34}"


"p=44"

d) if p=40, then Q=4


"Profit=TR-TC"


"TR=pQ=40\times4=160"


"Profit=160-170=-10"